there are 52 cards in a deck:
a.P(3 and black card)
b.P(3 or black card)

Answer by stanbon(75887)

there are 52 cards in a deck:
a.P(3 and black card) = 2/52
b.P(3 or black= P(3) + P(black) - P(3 and black)
= 4/52 + 26/52 - (2/52)
= 28/52
==========
Cheers,
Stan H.